OpenStudy (anonymous):

Find the second term in this sequence: tn=(6-n)^2

OpenStudy (anonymous):

6 36 16 9

OpenStudy (anonymous):

16

OpenStudy (anonymous):

t2 = (6-2)^2 = 4^2 = 16